# Exploring the Global-to-Local Attention Scheme in Graph Transformers: An Empirical Study

### v1

@@ 1 Introduction
Graph Neural Networks (GNNs), while serving as the dominant approach for graph representation learning, inherently suffer from over-smoothing and over-squashing issues alon2021on . Consequently, Graph Transformers (GTs) have recently gained research attention Chen22a , as their global attention mechanisms vaswani2017attention can mitigate these issues by enabling direct long-range dependency modeling without relying on stacked local aggregation. Nevertheless, recent analyses reveal that GTs may overemphasize distant neighbors, a limitation termed the over-globalizing problem pmlr-v235-xing24b .
@@
Integrating GNNs with the attention mechanism in GTs has become a prevalent approach, where GNNs model local information while the attention mechanism captures global dependencies ma2024polyformer ; 10479175 . There are two primary schemes for integrating GNNs and GTs as shown in Figure 1 (a) and (b): the local-and-global scheme where local layers and global layers output in parallel, and the local-to-global scheme where local layers precede global layers. Although both schemes demonstrate comparable performance with no consensus on superiority min2022transformergraphsoverviewarchitecture , and have even developed linear GTs wu2023sgformer ; deng2024polynormer with effective efficiency improvements, they still exhibit inherent limitations. First, in the local-and-global scheme, the integration method (summation or concatenation) is believed to be oversimplified and fails to facilitate effective information exchange ijcai2023p501 . Second, the local-to-global scheme is considered inferior due to inherent GNN limitations that can cause information loss and further propagation of noisy/erroneous features to downstream global layers rampasek2022GPS . Third, the over-globalizing tendency in GTs introduces structural imbalance: global attention in deeper layers exhibits a preference for distant node interactions, with comparatively insufficient utilization of local patterns from shallow layers, thereby compromising local feature extraction in the representation learning.
@@
  • To bridge the semantic gap between global attention and local GNN layers, we employ a novel cross-layer information fusion strategy that dynamically balances global and local information, which adaptively reallocates node weights, and effectively addresses both the over-smoothing problem in local layers and the over-globalization issue in attention layers. Remarkably, the entire fusion process does not increase model complexity, demonstrating that enhanced cross-layer interaction need not compromise computational efficiency.
  • We propose G2LFormer as an implementation of the global-to-local attention scheme and employ an existing attention mechanism (derived from SGFormer wu2023sgformer ) and a GNN (Cluster-GCN clustergcn or GatedGCN gatedgcn ) as the backbones. Experiments show that G2LFormer achieves state-of-the-art performance on both node-level tasks and graph-level tasks. We also prove the linear complexity of G2LFormer theoretically and experimentally.
@@ 2 Related Work
Graph Neural Networks. Message passing neural network (MPNN) is a general paradigm of GNNs, which utilizes message passing to propagate and aggregate information from local structures, enabling it to achieve excellent performance in graph representation learning pmlr-v70-gilmer17a . Numerous GNN advancements are derived from classic GNNs such as graph convolutional network (GCN) kipf2017semisupervisedclassificationgraphconvolutional and Graph Attention Network (GAT) velickovic2018graph , which are popular baselines for graph representation learning. Classic GNNs are the fundamental models for adapting graphs and have already undergone extensive algorithmic optimizations NEURIPS2022_1385753b . Currently, adapting to large-scale graphs is a challenge for graph mining models. For large-scale graph processing, GNN-based approaches are mainly aimed at optimizing training strategies 10.1145/3580305.3599565 ; 10.1007/978-3-031-26390-3_22 ; clustergcn . Cluster-GCN is such an improvement on GCN, incorporating graph partitioning and local convolution, tailored for large-scale graph adaptation clustergcn . To address the drawbacks of over-smoothing and over-squeezing of GNNs, GatedGCN gatedgcn augments standard graph convolutions with gating mechanisms to dynamically regulate message passing between nodes. Furthermore, to effectively transmit information within the network layers, a cross-layer information fusion strategy is proposed in wang2024nodescreatedequalnodespecific , which is employed in this work to selectively filter and propagate beneficial information between global layers and successive local layers.
@@
Graph Transformers. Early GTs suffered from limited scalability due to their quadratic complexity and exhibited relatively rudimentary expressivity shehzad2024graphtransformerssurvey . The high memory requirements of complex datasets highlight the importance of scalability, prompting a growing number of GTs to conduct experimental evaluations on complex datasets kong2023goat ; zhang2024torchgtholisticlargescalegraph . To balance scalability and expressivity, algorithm-level optimizations are prevalent. Some GTs tend to exhibit linear complexity while maintaining performance, and run efficiently on large-scale graphs ma2024polyformer ; yang2022hypformer . However, prioritizing running speed alone is not the optimal solution, as overly concise propagation formulas are not conducive to the model learning complex node representations Li2024Rethinking . Moreover, few researchers have yet discussed the categories of attention schemes at the model-level. Despite being regarded as having limitations, the local-to-global scheme and the local-and-global scheme remain two primary attention schemes ijcai2023p501 . The local-to-global attention scheme is the most commonly employed architecture in early GTs, represented by GraphTrans NEURIPS2021_6e67691b . The local-and-global attention scheme subsequently garners increased scrutiny as the sequential transmission between local layers and global layers is believed to cause information loss, with GraphGPS being a notable representative of this scheme rampasek2022GPS . Nevertheless, the approach of parallel output between local layers and global layers within the local-and-global attention scheme is criticized for being simplistic ijcai2023p501 . At present, the representative GTs for the local-and-global attention scheme and the local-to-global attention scheme are SGFormer wu2023sgformer and Polynormer deng2024polynormer , respectively. SGFormer proposes a linear attention mechanism that obviates approximation processing, demonstrating high representational capacity with a single attention network layer. Polynormer employs polynomial functions in Transformer’s formula, streamlining the attention mechanism while exhibiting superior performance. The aforementioned two models make a comprehensive trade-off between expressivity and scalability. The attention mechanism of SGFormer has extremely low overhead and is employed as our backbone for global layers.
@@ 3 Preliminaries
Attention mechanism. GTs build upon the Transformer architecture vaswani2017attention , whose core operation is computing self-attention. The standard self-attention mechanism consists of three learnable matrices, i.e. the query (Q), key (K) and value (V) matrices. The three matrices are calculated based on node features $X$ and specific weights as in Equation 2, where $W_{Q}$, $W_{K}$, $W_{V}$ $\in\mathbb{R}^{d\times d}$ are the weight matrices corresponding to $Q$, $K$, and $V$.
@@ 4.1 Motivation
Both local-and-global and local-to-global places global layers near the deep parts of networks, resulting in global-layer-dominated learning of final node representations. According to recent evaluations on GTs, the over-globalizing problem prevents nodes from sharing the same label with their proximate neighbors pmlr-v235-xing24b , which means that global layers underutilize local information. Thus, a proposal is put forward: to adopt a global-to-local attention scheme, i.e., positioning global layers in the shallow parts while placing local layers in the deep part, thereby alleviating the deficiencies of the attention mechanism.
@@ 4.2 Architecture
As a concrete implementation of the global-to-local scheme, we propose G2LFormer, a versatile model adaptable to diverse downstream tasks that employs Cluster-GCN clustergcn or GatedGCN gatedgcn as the backbone of local layers and the global attention component of SGFormer wu2023sgformer as the backbone of global layers.
@@ 4.3.2 Cross-layer information fusion strategy
To preserve valuable information from $h_{TL}$ in local layers, the cross-layer information fusion strategy is introduced before local layers. Specifically, G2LFormer employs NOSAF (Node-Specific Layer Aggregation and Filtration) NEURIPS2022_092359ce ; wang2024nodescreatedequalnodespecific to ensure effective information flow to the next layer, thereby preventing over-smoothing in local layers, while simultaneously reallocating node weights to preserve critical information of global layers and mitigate over-globalization. NOSAF redesigns node aggregation, including the refinement of node weights, which is applicable to improve weight allocation in attention mechanisms. The process of reallocating node weight is as follows:
@@ 5.1 Datasets
For node classification tasks, we select large-scale graphs commonly used in the literature. The social network dataset Pokec Leskovec_Krevl_2014 is chosen for its characteristics as a large-scale heterogeneous graph. From the Open Graph Benchmark (OGB) hu2020ogb , we consider three representative graphs: ogbn-arxiv, ogbn-proteins, and ogbn-products, which cover different domains and scales. ogbn-products, also known as amazon2m, is an Amazon product co-purchasing network. ogbn-proteins is a protein interaction network. ogbn-arxiv, a citation network comprising 0.16 million nodes and 1.1 million edges, is the only dataset in this study that permits full-batch training. Specifically, for ogbn-arxiv, we adopt the TAPE framework he2024harnessingexplanationsllmtolminterpreter and use the node features it provides. The reason is that TAPE employs a large language model to enrich the node features of the text attribute graphs, which improves the training speed by fine-tuning the original texts through the language model, and achieves state-of-the-art performance in ogbn-arxiv.
@@ 5.2 Baselines
For node classification tasks, we compare our method with classic GNN baselines (GCN and GAT velickovic2018graph ), which are widely adopted due to their moderate complexity and satisfactory expressivity. Furthermore, SIGN frasca2020signscalableinceptiongraph and SGC pmlr-v97-wu19e are considered, as they are GNNs that focus on improving scalability. Cluster-GCN clustergcn is not included in our results, as it shows negligible accuracy difference from baseline GCN. Among various GT variants, DIFFormer wu2023difformer , SGFormer wu2023sgformer , and Polynormer deng2024polynormer emerge as notable representatives. These linear GTs offer not only straightforward implementation but also enhanced expressivity and scalability.
For graph-level tasks, we adopt three classic GNN baselines: GCN, GIN xu2019powerfulgraphneuralnetworks , and GatedGCN. As competing graph transformer baselines, we select GraphGPS rampasek2022GPS , Exphormer shirzad2023exphormer , and GRIT ma2023GraphInductiveBiases based on their demonstrated superiority in graph-level tasks. We also include GECO sancak2024scalableeffectivealternativegraph , an enhanced model that replaces the attention mechanism, representing a novel variant of GTs. To capture long-range dependencies, we additionally evaluate state space models through two representative implementations: GPS+Mamba Behrouz2024KDD-GMN and GMN Behrouz2024KDD-GMN .
@@ 5.3 Setup
The data splitting for each dataset is based on existing practices. Specifically, ogbn-arxiv, ogbn-proteins, and ogbg-molhiv are split according to OGB standard settings. All datasets from LRGB are split following the standard protocol. ogbn-arxiv is chronologically partitioned, with papers published until 2017 allocated for training, those published in 2018 for validation, and papers from 2019 reserved for testing. ogbn-proteins is split based on the species of proteins. ogbn-products and pokec datasets are processed following recent work wu2022nodeformer ; wu2023difformer ; wu2023sgformer . The data splitting ratio of ogbn-products is 50% / 25% / 25% (train / val / test), while pokec follows a 10% / 10% / 80% split.
@@ 5.5 Ablation study
In this work, we employ the cross-layer information fusion strategy in the implementation of a global-to-local attention scheme. To systematically evaluate the contributions of each component, we conduct ablation studies examining (1) the attention scheme variants and (2) the cross-layer fusion strategy independently. Although the local-and-global attention scheme is theoretically independent of cross-layer information fusion strategies, empirical evidence suggests that such strategies can enhance its effectiveness. Therefore, we empirically validate the influence of the cross-layer information fusion strategy on the local-and-global attention scheme, as well. In summary, the experimental modules encompass three attention schemes: global-to-local scheme, local-to-global scheme, and local-and-global scheme, with or without the cross-layer information fusion strategy. The performance of each module is evaluated on pokec, with results presented in Figure 4.
@@
Figure 4 indicates that the cross-layer information fusion strategy employed in this work benefits both the global-to-local and local-to-global attention schemes, whereas it does not enhance the performance of the local-and-global attention scheme, which implies that the sequential output methodology between global layers and local layers is susceptible to information loss and requires corresponding measures to prevent it. Meanwhile, the global-to-local scheme incorporating a cross-layer information fusion strategy consistently maintains optimal performance, which is likely attributable to the alleviation of the over-globalizing problem in global layers.
Our findings also align with previous research suggesting that existing attention schemes exhibit inherent limitations ijcai2023p501 , and the challenges associated with GTs underscore the exploration of novel attention schemes pmlr-v235-xing24b . Note that this work primarily focuses on demonstrating the feasibility of the global-to-local attention scheme. Its optimality remains to be fully investigated, as the expressiveness of attention mechanisms is typically backbone-model-dependent.
@@ 5.6 Efficiency
To evaluate the efficiency of G2LFormer, we follow the approach of FairGT fairgt2024luo . Table 3 presents the per-epoch training time and GPU memory usage of G2LFormer compared with representative baselines on the ogbn-proteins and pokec datasets. When considering the performance advantages demonstrated in the experimental results section, G2LFormer has achieved a favorable balance between expressivity and computational cost. Although it incurs higher training time than SGFormer and GCN variants due to the additional cross-layer information fusion strategy, it consistently consumes less GPU memory (18.4-45.2% lower) than Polynormer. Notably, Polynormer runs out of memory (OOM) on ogbn-proteins when the batch size reaches 50,000, whereas G2LFormer remains stable. Moreover, the G2LFormer backbone can be further optimized or replaced with more efficient models to accelerate training in future work.

### v2

@@ 1 Introduction
Graph Neural Networks (GNNs), while serving as the dominant approach for graph representation learning, inherently suffer from over-smoothing and over-squashing issues Alon and Yahav (2021). Consequently, Graph Transformers (GTs) have recently gained research attention Chen et al. (2022), as their global attention mechanisms Vaswani et al. (2017) can mitigate these issues by enabling direct long-range dependency modeling without relying on stacked local aggregation. Nevertheless, recent analyses reveal that GTs may overemphasize distant neighbors, a limitation termed the over-globalizing problem Xing et al. (2024).
@@
Integrating GNNs with the attention mechanism in GTs has become a prevalent approach, where GNNs model local information while the attention mechanism captures global dependencies Ma et al. (2024); Kuang et al. (2024). There are two primary schemes for integrating GNNs and GTs as shown in Figure 1 (a) and (b): the local-and-global scheme where local layers and global layers output in parallel, and the local-to-global scheme where local layers precede global layers. Although both schemes demonstrate comparable performance with no consensus on superiority Min et al. (2022), and have even developed linear GTs Wu et al. (2023b); Deng et al. (2024) with effective efficiency improvements, they still exhibit inherent limitations. First, in the local-and-global scheme, the integration method (summation or concatenation) is believed to be oversimplified and fails to facilitate effective information exchange Yin and Zhong (2023). Second, the local-to-global scheme is considered inferior due to inherent GNN limitations that can cause information loss and further propagation of noisy/erroneous features to downstream global layers Rampášek et al. (2022). Third, the over-globalizing tendency in GTs introduces structural imbalance: global attention in deeper layers exhibits a preference for distant node interactions, with comparatively insufficient utilization of local patterns from shallow layers, thereby compromising local feature extraction in the representation learning.
@@
  • To prevent information loss between network layers, we employ a novel cross-layer information fusion strategy that dynamically balances global and local information, which adaptively reallocates node weights, and effectively addresses both the over-smoothing problem in local layers and the over-globalization issue in attention layers. Remarkably, the entire fusion process does not increase model complexity, demonstrating that enhanced cross-layer interaction need not compromise computational efficiency.
  • We propose G2LFormer as an implementation of the global-to-local attention scheme and employ an existing attention mechanism (derived from SGFormer Wu et al. (2023b)) and a GNN (Cluster-GCN Chiang et al. (2019) or GatedGCN Bresson and Laurent (2017)) as the backbones. Experiments show that G2LFormer achieves state-of-the-art performance on both node-level tasks and graph-level tasks. We also prove the linear complexity of G2LFormer theoretically and experimentally.
@@ 2 Related Work
Graph Neural Networks. Message passing neural network (MPNN) is a general paradigm of GNNs, which utilizes message passing to propagate and aggregate information from local structures, enabling it to achieve excellent performance in graph representation learning Gilmer et al. (2017). Numerous GNN advancements are derived from classic GNNs such as graph convolutional network (GCN) Kipf and Welling (2017) and Graph Attention Network (GAT) Veličković et al. (2018), which are popular baselines for graph representation learning. Classic GNNs are the fundamental models for adapting graphs and have already undergone extensive algorithmic optimizations Ding et al. (2022). Currently, adapting to large-scale graphs is a challenge for graph mining models. For large-scale graph processing, GNN-based approaches are mainly aimed at optimizing training strategies Xue et al. (2023); Maekawa et al. (2023); Chiang et al. (2019). Cluster-GCN is such an improvement on GCN, incorporating graph partitioning and local convolution, tailored for large-scale graph adaptation Chiang et al. (2019). To address the drawbacks of over-smoothing and over-squeezing of GNNs, GatedGCN Bresson and Laurent (2017) augments standard graph convolutions with gating mechanisms to dynamically regulate message passing between nodes. Furthermore, to effectively transmit information within the network layers, a cross-layer information fusion strategy is proposed in Wang et al. (2024), which is employed in this work to selectively filter and propagate beneficial information between global layers and successive local layers.
@@
Graph Transformers. Early GTs suffered from limited scalability due to their quadratic complexity and exhibited relatively rudimentary expressivity Shehzad et al. (2024). The high memory requirements of complex datasets highlight the importance of scalability, prompting a growing number of GTs to conduct experimental evaluations on complex datasets Kong et al. (2023); Zhang et al. (2024). To balance scalability and expressivity, algorithm-level optimizations are prevalent. Some GTs tend to exhibit linear complexity while maintaining performance, and run efficiently on large-scale graphs Ma et al. (2024); Yang et al. (2024). However, prioritizing running speed alone is not the optimal solution, as overly concise propagation formulas are not conducive to the model learning complex node representations Li et al. (2024). Moreover, few researchers have yet discussed the categories of attention schemes at the model-level. Despite being regarded as having limitations, the local-to-global scheme and the local-and-global scheme remain two primary attention schemes Yin and Zhong (2023). The local-to-global attention scheme is the most commonly employed architecture in early GTs, represented by GraphTrans Wu et al. (2021). The local-and-global attention scheme subsequently garners increased scrutiny as the sequential transmission between local layers and global layers is believed to cause information loss, with GraphGPS being a notable representative of this scheme Rampášek et al. (2022). Nevertheless, the approach of parallel output between local layers and global layers within the local-and-global attention scheme is criticized for being simplistic Yin and Zhong (2023). At present, the representative GTs for the local-and-global attention scheme and the local-to-global attention scheme are SGFormer Wu et al. (2023b) and Polynormer Deng et al. (2024), respectively. SGFormer proposes a linear attention mechanism that obviates approximation processing, demonstrating high representational capacity with a single attention network layer. Polynormer employs polynomial functions in Transformer’s formula, streamlining the attention mechanism while exhibiting superior performance. The aforementioned two models make a comprehensive trade-off between expressivity and scalability. The attention mechanism of SGFormer has extremely low overhead and is employed as our backbone for global layers.
@@ 3 Preliminaries
Attention mechanism. GTs build upon the Transformer architecture Vaswani et al. (2017), whose core operation is computing self-attention. The standard self-attention mechanism consists of three learnable matrices, i.e. the query (Q), key (K) and value (V) matrices. The three matrices are calculated based on node features $X$ and specific weights as in Equation 2, where $W_{Q}$, $W_{K}$, $W_{V}$ $\in\mathbb{R}^{d\times d}$ are the weight matrices corresponding to $Q$, $K$, and $V$.
@@ 4.1 Motivation
Both local-and-global and local-to-global places global layers near the deep parts of networks, resulting in global-layer-dominated learning of final node representations. According to recent evaluations on GTs, the over-globalizing problem prevents nodes from sharing the same label with their proximate neighbors Xing et al. (2024), which means that global layers underutilize local information. Thus, a proposal is put forward: to adopt a global-to-local attention scheme, i.e., positioning global layers in the shallow parts while placing local layers in the deep part, thereby alleviating the deficiencies of the attention mechanism.
@@ 4.2 Architecture
As a concrete implementation of the global-to-local scheme, we propose G2LFormer, a versatile model adaptable to diverse downstream tasks that employs Cluster-GCN Chiang et al. (2019) or GatedGCN Bresson and Laurent (2017) as the backbone of local layers and the global attention component of SGFormer Wu et al. (2023b) as the backbone of global layers.
@@ 4.3.2 Cross-layer information fusion strategy
To preserve valuable information from $h_{TL}$ in local layers, the cross-layer information fusion strategy is introduced before local layers. Specifically, G2LFormer employs NOSAF (Node-Specific Layer Aggregation and Filtration) Luan et al. (2022); Wang et al. (2024) to ensure effective information flow to the next layer, thereby preventing over-smoothing in local layers, while simultaneously reallocating node weights to preserve critical information of global layers and mitigate over-globalization. NOSAF redesigns node aggregation, including the refinement of node weights, which is applicable to improve weight allocation in attention mechanisms. The process of reallocating node weight is as follows:
@@ 5.1 Datasets
For node classification tasks, we select large-scale graphs commonly used in the literature. The social network dataset Pokec Leskovec and Krevl (2014) is chosen for its characteristics as a large-scale heterogeneous graph. From the Open Graph Benchmark (OGB) Hu et al. (2020), we consider three representative graphs: ogbn-arxiv, ogbn-proteins, and ogbn-products, which cover different domains and scales. ogbn-products, also known as amazon2m, is an Amazon product co-purchasing network. ogbn-proteins is a protein interaction network. ogbn-arxiv, a citation network comprising 0.16 million nodes and 1.1 million edges, is the only dataset in this study that permits full-batch training. Specifically, for ogbn-arxiv, we adopt the TAPE framework He et al. (2024) and use the node features it provides. The reason is that TAPE employs a large language model to enrich the node features of the text attribute graphs, which improves the training speed by fine-tuning the original texts through the language model, and achieves state-of-the-art performance in ogbn-arxiv.
@@ 5.2 Baselines
For node classification tasks, we compare our method with classic GNN baselines (GCN and GAT Veličković et al. (2018)), which are widely adopted due to their moderate complexity and satisfactory expressivity. Furthermore, SIGN Frasca et al. (2020) and SGC Wu et al. (2019) are considered, as they are GNNs that focus on improving scalability. Cluster-GCN Chiang et al. (2019) is not included in our results, as it shows negligible accuracy difference from baseline GCN. Among various GT variants, DIFFormer Wu et al. (2023a), SGFormer Wu et al. (2023b), and Polynormer Deng et al. (2024) emerge as notable representatives. These linear GTs offer not only straightforward implementation but also enhanced expressivity and scalability.
For graph-level tasks, we adopt three classic GNN baselines: GCN, GIN Xu et al. (2019), and GatedGCN. As competing graph transformer baselines, we select GraphGPS Rampášek et al. (2022), Exphormer Shirzad et al. (2023), and GRIT Ma et al. (2023) based on their demonstrated superiority in graph-level tasks. We also include GECO Sancak et al. (2024), an enhanced model that replaces the attention mechanism, representing a novel variant of GTs. To capture long-range dependencies, we additionally evaluate state space models through two representative implementations: GPS+Mamba Behrouz and Hashemi (2024) and GMN Behrouz and Hashemi (2024).
@@ 5.3 Setup
The data splitting for each dataset is based on existing practices. Specifically, ogbn-arxiv, ogbn-proteins, and ogbg-molhiv are split according to OGB standard settings. All datasets from LRGB are split following the standard protocol. ogbn-arxiv is chronologically partitioned, with papers published until 2017 allocated for training, those published in 2018 for validation, and papers from 2019 reserved for testing. ogbn-proteins is split based on the species of proteins. ogbn-products and pokec datasets are processed following recent work Wu et al. (2022, 2023a, 2023b). The data splitting ratio of ogbn-products is 50% / 25% / 25% (train / val / test), while pokec follows a 10% / 10% / 80% split.
@@ 5.5 Ablation study
In this work, we employ the cross-layer information fusion strategy in the implementation of a global-to-local attention scheme. To systematically evaluate the contributions of each component, we conduct ablation studies examining (1) the attention scheme variants and (2) the cross-layer fusion strategy independently. Although the local-and-global attention scheme is theoretically independent of cross-layer information fusion strategies, empirical evidence suggests that such strategies can enhance its effectiveness. Therefore, we empirically validate the influence of the cross-layer information fusion strategy on the local-and-global attention scheme, as well. In summary, the experimental modules encompass three attention schemes: global-to-local scheme, local-to-global scheme, and local-and-global scheme, with or without the cross-layer information fusion strategy. The performance of each module is evaluated on Peptides-struct and pokec, with results presented in Figure 4.
@@
Figure 4 indicates that the cross-layer information fusion strategy employed in this work benefits the global-to-local and local-to-global attention schemes, which implies that the sequential output methodology between global layers and local layers is susceptible to information loss and requires corresponding measures to prevent it. Since both local-to-global and local-and-global attention schemes are suboptimal and incorporate more misleading semantic information, cross-layer information fusion strategies may fail to yield consistent performance improvements. Meanwhile, the global-to-local scheme incorporating a cross-layer information fusion strategy consistently maintains optimal performance, which is likely attributable to the alleviation of the over-globalizing problem in global layers.
Our findings also align with previous research suggesting that existing attention schemes exhibit inherent limitations Yin and Zhong (2023), and the challenges associated with GTs underscore the exploration of novel attention schemes Xing et al. (2024). Note that this work primarily focuses on demonstrating the feasibility of the global-to-local attention scheme. Its optimality remains to be fully investigated, as the expressiveness of attention mechanisms is typically backbone-model-dependent.
@@ 5.6 Efficiency
To evaluate the efficiency of G2LFormer, we follow the approach of FairGT Luo et al. (2024). Table 3 presents the per-epoch training time and GPU memory usage of G2LFormer compared with representative baselines on the ogbn-proteins and pokec datasets. When considering the performance advantages demonstrated in the experimental results section, G2LFormer has achieved a favorable balance between expressivity and computational cost. Although it incurs higher training time than SGFormer and GCN variants due to the additional cross-layer information fusion strategy, it consistently consumes less GPU memory (18.4-45.2% lower) than Polynormer. Notably, Polynormer runs out of memory (OOM) on ogbn-proteins when the batch size reaches 50,000, whereas G2LFormer remains stable. Moreover, the G2LFormer backbone can be further optimized or replaced with more efficient models to accelerate training in future work.